The demand function and the average cost .Determine the:

Profit function.

Profit. Hence show it is maximum


1
Expert's answer
2022-03-07T11:42:31-0500

Demand function shows the functional relationship between quantity demanded for a commodity and its various determinants. Quantity demanded is always inversely related to price. The formula for a demand function is Q=a + bP, where a is factors affecting demand except price, b is the slope of the curve, and P is the price. To get the profit function, an inverse of demand function is done. Therefore, if the demand function is Q=240-2P, then the inverse demand function would be P=120-0.5Q. This makes the Profit function to be P=120-0.5Q.

To get maximum profit, we find where marginal revenue is equal to marginal profit by looking at the point of intersection. The point of intersection shows maximum profit of the company.